OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

plzz help

OpenStudy (isaiah.feynman):

Alright first thing to remember is perpendicular lines have a slope m1m2=-1. We have slope 2 for the first line the slope for the second line will be -1/2. Okay now we have the slope for the second line, next thing is to get the equation. To do this we use the points (6,3) we slot in 6 and 3 for x1 and y1 in the equation y-y1= m(x-x1) (also remember we already have m which is -1/2!!!) slotting in gives us y-3=-1/2(x-6). Simplifying the equation gives y=-1/2x+6!!! And that's your equation!!! :)

OpenStudy (anonymous):

thankyou